# A Copula-Based Variational Autoencoder for Uncertainty Quantification in Inverse Problems: Application to Damage Identification in an Offshore Wind Turbine

### v1

@@ 2.4.2 Gaussian Copula
Copulas are functions that join multivariate distribution functions to their one-dimensional marginal distribution functions. According to Sklar’s theorem[33], any multivariate joint distribution can be represented by its marginal distributions and a corresponding copula, which uniquely describes the dependence structure between the variables. This decomposition is fundamental to their application, as it permits the separate modeling of the marginal behaviors and their interdependence. Such flexibility is particularly advantageous for constructing complex multivariate models. For a foundational treatment of copula theory, see Nelsen [64], and for an extensive survey of dependence modeling applications, see Joe [32]. Further mathematical details are provided in Appendix A.
@@ 3 Neural Network design and training
Regarding the encoder ($\mathcal{I}_{\bm{\theta}}$), the hidden layers keep the same architecture as [29]. At the output layer of the encoder, we consider different activation functions depending on the estimated features. We use the sigmoid function [75] for the means vector $\bm{\mu}$, as it yields a smooth output in the interval $[0,1]$. We use a scaled sigmoid activation for the rotation angles $\bm{\gamma}$, constraining them to the interval $0,2\pi]$, which corresponds to the domain of angles in radians. For the variances and scaling values, we use the softplus function [76], which is a smooth equivalent of ReLU that enforces positive values. For the Copula approach, we must estimate a lower triangular matrix, where we employ linear activation for the off-diagonal entries ($\mathbf{l}_{\text{off}}$), and softplus for the diagonal entries ($\mathbf{l}_{\text{diag}}$). Finally, we use softmax function [12] for the weights $\bm{\alpha}$, ensuring that their sum is equal to one and each value ranges into $[0,1]$. We denote with $\mathbf{p}_{\bm{\theta}}\in\mathbb{R}^{n_{p}}$ the vector of properties to be estimated by the encoder. The total number of properties $n_{p}$ depends on the chosen distributional model. We assume the same value for the noise parameter leveraging the contribution of both loss terms $\beta=0.075$, according to our previous analysis in [77]. Table 1 summarizes the architecture details of the encoder and decoder.
@@
In the latent space, we incorporate a sampling layer that builds the posterior distribution $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ from the estimated properties $\mathbf{p}_{\bm{\theta}}$ and draws $H$ samples within the latent space domain. The specifications of the sampling layer also depend on the selected distributional model (see the details in sections 2.4.1 and 2.4.2). The drawn samples enter the decoder together with the observed operating conditions in $\mathbf{w}$. The number of samples to draw is set to $H=1$ as it suffices to ensure convergence. A larger number of samples implies more time-consuming iterations. Following the experience of the authors according to previous work [77], we assume a noise parameter $\beta=0.075$, corresponding (roughly) to noise levels of up to $8.5\%$ [78]. This parameter affects the contribution of the distributional term in Eq. 19.
@@ 4.2 Performance analysis
where $\text{LL}_{i}$ denotes the model log likelihood, $n$ indicates the number of points in the cloud to be fitted by the model ($N^{f}$ in Eqs. 32 and 33), and $p_{i}$ is the number of parameters in the model. The first term, common to both expressions, reflects how well the model fits the data according to the log-likelihood metric, which decreases as the fitness (likelihood) increases. This term often governs the overall trend of the evolution of both metrics. The second terms penalize the number of parameters in the model.

### v2

@@ 2.4.2 Gaussian Copula
Herein, we propose a Copula-based distributional model [53] to tackle the scalability limitations of full-covariance GMs. Copulas are functions that join the multivariate distribution (describing the interrelation among variables) to the independent one-dimensional marginal distributions in an efficient manner [54, 55]. According to Sklar’s theorem[90], any multivariate joint distribution can be represented by its marginal distributions and a corresponding copula, which uniquely describes the dependence structure between the variables. This decomposition is fundamental to their application, as it permits the separate modeling of the marginal behaviors and their interdependence. Such flexibility is particularly advantageous for constructing complex multivariate models. Copulas can adopt various forms, mainly divided into (a) parametric (e.g., Gaussian, Archimedean), or (b) non-parametric (e.g., Clayton, Gumbel) [91, 92]. For a foundational treatment of copula theory, see Nelsen [93], and for an extensive survey of dependence modeling applications, see Joe [55]. Further mathematical details are provided in Appendix A.
@@ 3 Neural Network design and training
Regarding the encoder ($\mathcal{I}_{\bm{\theta}}$), the hidden layers keep the same architecture as [52]. At the output layer of the encoder, we consider different activation functions depending on the estimated features. We use the sigmoid function [104] for the means vector $\bm{\mu}$, as it yields a smooth output in the interval $[0,1]$. We use a scaled sigmoid activation for the rotation angles $\bm{\gamma}$, constraining them to the interval $0,2\pi]$, which corresponds to the domain of angles in radians. For the variances and scaling values, we use the softplus function [105], which is a smooth equivalent of ReLU that enforces positive values. For the Copula approach, we must estimate a lower triangular matrix, where we employ linear activation for the off-diagonal entries ($\mathbf{l}_{\text{off}}$), and softplus for the diagonal entries ($\mathbf{l}_{\text{diag}}$). Finally, we use softmax function [13] for the weights $\bm{\alpha}$, ensuring that their sum is equal to one and each value ranges into $[0,1]$. We denote with $\bm{\zeta}_{\bm{\theta}}\in\mathbb{R}^{n_{p}}$ the vector of properties to be estimated by the encoder. The total number of properties $n_{p}$ depends on the chosen distributional model. We assume the same value for the noise parameter leveraging the contribution of both loss terms $\beta=0.075$, according to our previous analysis in [106]. Table 1 summarizes the architecture details of the encoder and decoder.
@@
In the latent space, we incorporate a sampling layer that builds the posterior distribution $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ from the estimated properties $\bm{\zeta}_{\bm{\theta}}$ and draws $H$ samples within the latent space domain. The specifications of the sampling layer also depend on the selected distributional model (see the details in sections 2.4.1 and 2.4.2). The drawn samples enter the decoder together with the observed operating conditions in $\mathbf{w}$. The number of samples to draw is set to $H=1$ as it suffices to ensure convergence. A larger number of samples implies more time-consuming iterations. Following the experience of the authors according to previous work [106], we assume a noise parameter $\beta=0.075$, corresponding (roughly) to noise levels of up to $8.5\%$ [107]. This parameter affects the contribution of the distributional term in eq. (19).
@@ 4.2 Performance analysis
where $\text{LL}_{i}$ denotes the model log likelihood, $n$ indicates the number of points in the cloud to be fitted by the model ($N^{f}$ in Eqs. (34) and (35)), and $p_{i}$ is the number of parameters in the model. The first term, common to both expressions, reflects how well the model fits the data according to the log-likelihood metric, which decreases as the fitness (likelihood) increases. This term often governs the overall trend of the evolution of both metrics. The second terms penalize the number of parameters in the model.